OpenStudy (anonymous):

HELP PLEASE WITH ALGEBRA QUESTION Rewrite in simplest rational exponent form square root x times 4 square root x show each step of your process

OpenStudy (anonymous):

@LexiLuvv2431 @Ashy98 @ali2x2 @bubblysmiles24 @BrandonR @babylove2015 @Crazyandbeautiful @ChiakiNagoya @Calcmathlete

OpenStudy (anonymous):

Is your question \(\sqrt{x}\times4\sqrt{x}\)?

OpenStudy (anonymous):

\[\sqrt{x} \times \sqrt[4]{x}\]

OpenStudy (anonymous):

i used to live in FM how crazy!

OpenStudy (anonymous):

oops sorry i asked the wrong question can i give you a new one? sorry i already figured this one out

OpenStudy (anonymous):

really!? and np :)

OpenStudy (anonymous):

This is the new question: Rewrite in simplest radical form: \[x^{5/6}\] divided by \[x^{1/6}\] Show each step of your process

OpenStudy (anonymous):

Alright. Do you remember what you do to the exponents when you divide with the same base?

OpenStudy (anonymous):

so it would be exponent 2/3 and then move the 3 outside the square root and the 2 inside?

OpenStudy (anonymous):

Yes! Excellent job!

OpenStudy (anonymous):

So, you should get \(\sqrt[3]{x^2}\) to make sure you're thinking what I'm thinking.

OpenStudy (anonymous):

And yeah, if you want a quick way to remember those exponent rules, just do a small calculation to make sure. The reason \(\frac{2^3}{2^2}=2^1\), for example, is because exponents are just multiplication. So, \frac{2\times2\times2}{2\times2}, and you can cancel two of those 2's, and you're left with one 2.

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

Sorry about the formatting, \[\frac{2\times2\times2}{2\times2}\]but I think you get the gist of it. And yeah, it's cool meeting someone else from FM. Gtg, see ya. :)

OpenStudy (anonymous):

its fine and yeah its funny! thanks again bye!